Let π‘₯π‘˜ be a Cauchy sequence and 𝑓 be a continuous function, show that 𝑓(π‘₯π‘˜) is also a Cauchy sequence.Β 


1
Expert's answer
2021-12-20T16:47:32-0500

Solution:

It is false.

Example:

Take

"\begin{gathered}

X=(0,+\infty) \\

x_{k}=\frac{1}{k} \\

f: x \mapsto \frac{1}{x}

\end{gathered}"

"\left(x_{k}\right)" is Cauchy

f is continuous

"f\left(x_{k}\right)=k" is not Cauchy

Given statement will be true if f is UNIFORMLY CONTINUOUS.